Ann has $7000.00 invested in two accounts, part at 7% simple interest and part at 5.25% simple interest for a period of 1 year. How much did she invest at each rate if each account earned the same interest?
Answer by mananth(16946)

Part I -----------x--------------7%
Part II----------y---------------5.25%
x+y=7000
7%x=5.25%y
7x=5.25y
7x-5.25y=0
1 x + 1 y = 7000 .............1
Total value
7 x -5 1/4 y = 0 .............2
Eliminate y
multiply (1)by 5 1/4
Multiply (2) by 1
5 1/4 x 5 1/4 y = 36750
7 x -5 1/4 y = 0
Add the two equations
12 1/4 x = 36750
/ 12 1/4
x = 3000
plug value of x in (1)
1 x + 1 y = 7000
3000 + y = 7000
y = 7000 -3000
y = 4000
y = 4000